OpenStudy (anonymous):

Find the range if y = 3x + 12 for x = {0.4, 0.6, 1.8, 2.2}.

OpenStudy (anonymous):

A. {1.2, 1.8, 5.4, 6.6} B. {13.2, 13.8, 17.4, 18.6} C. {12.4, 12.6, 13.8, 14.2} Idk whitch one it is I'm stuck

OpenStudy (ivettef365):

just substitute the x for (0,4,0 .......) and see what is the value of y that is the range

OpenStudy (ivettef365):

example when x = 0.4 then y = 3x+12 y = 3(0.4) + 12 y= 1.20 + 12 = 13.20

OpenStudy (ivettef365):

now find out what is y when x = 0.6